# EquiContact: A Hierarchical SE(3) Vision-to-Force Equivariant Policy for Spatially Generalizable Contact-rich Tasks

### v1

@@ IV-B2 G-CompACT
where $g=(p,R)\in SE(3)$ is a current end-effector pose and $g_{{}_{EDF}}=(p_{{}_{EDF}},R_{{}_{EDF}})$ is the reference frame predicted by Diff-EDF. The derivation of the GCEV $e_{{}_{G}}$ is presented in [12]. We note that the GCEV denotes an error vector on the end-effector body frame as emphasized in [11].
The model outputs a chunk of $N$ relative pose actions:
$\bar{g}_{d}(k+i)=g(k)\cdot g_{rel}(k+i),\quad i=1,\cdots,N,$ (7)
where each $g_{rel}(k+i)\in SE(3)$ is the $i$th relative pose from the ACT model and $\bar{g}_{d}(k+i)$ is the $i$th temporary desired pose signal. A temporal ensembling mechanism is then applied to $\bar{g}_{d}(k+i)$ to compute the desired pose $g_{d}(k)$ for GAC. To respect the manifold structure of $SE(3)$, we use the scipy implementation for weighted averaging on the rotation matrix to get $R_{d}(k)$ and standard weighted averaging on positions for $p_{d}(k)$ separately.
Our proposed G-CompACT also outputs the admittance gains matrices $K_{p}$ and $K_{R}$ in the GAC elastic wrench in Eq. (4). Unlike CompACT, the gain matrices are represented in the desired frame of the end-effector [13], removing the necessity of utilizing full-gain matrix representations and coordinate transformations to deal with spatial equivariance. Therefore, a diagonal matrix representation in $K_{p}$ and $K_{R}$ is sufficient to conduct the PiH task successfully, without requiring a Cholesky factorization to represent the gains, as in CompACT. Since the diagonal component representation of matrices is equivalent to a vector, the standard temporal ensemble can be applied. Similar to the relative pose action, the ACT provides temporary gain signals $\bar{K}_{p}$ and $\bar{K}_{R}$, followed by a temporal ensemble to obtain $K_{p}$ and $K_{R}$.
Formally, let the G-CompACT method be $\pi_{\theta}$ such that
$\begin{split}a(k)&=\pi_{\theta_{2}}(o(k)),\quad\text{where}\\ a(k)&\triangleq\{(g_{rel}(k+i),\bar{K}_{p}(k+i),\bar{K}_{R}(k+i))\}_{i=1}^{N},\\ o(k)&\triangleq(e_{{}_{G}},F_{e},I_{w})(k),\end{split}$ (8)
where $\bar{K}_{p}$ and $\bar{K}_{R}$ denote the temporary admittance gains, $a(k)$ denotes the actions and $o(k)$ denotes the observation at time step $k$, $e_{{}_{G}}$ is GCEV defined in (6), $F_{e}$ is force-torque sensor reading on the end-effector frame, and $I_{w}=\{I_{w,1},I_{w,2}\}$ where $I_{w,1}$ and $I_{w,2}$ denote the images from the wrist cameras #1 and #2, respectively (Fig. 2).
@@
Figure 3: Effects of the left group action $g_{l}$ to the end-effector pose $g$ and the reference frame $g_{ref}$, and to the wrists cameras $I_{w,1}$ and $I_{w,2}$. As the left group action is applied on the end-effector, the wrist cameras start to see not only the optical table but also the outer backgrounds. $\{s\}$ denotes a spatial frame.
@@
The left-translated observation signals $o(k)$ reads that:
$g_{l}\circ o(k)=(g_{l}\circ e_{{}_{G}},g_{l}\circ F_{e},g_{l}\circ I_{w}).$ (12)
As was shown in Lemma 1 of [11], the GCEV $e_{{}_{G}}$ is left invariant as
$g_{l}\circ e_{{}_{G}}(g,g_{{}_{EDF}})=e_{{}_{G}}(g_{l}g,g_{l}g_{{}_{EDF}})=e_{{}_{G}}(g,g_{{}_{EDF}}).$
The force-torque sensor values are left-invariant because they are already defined with respect to the end-effector frame [11], and the visual representation vectors satisfy left invariance due to Assumption 1. Combining all these properties, it follows that
$a(k)=\pi_{\theta_{2}}(g_{l}\circ o(k))=\pi^{\prime}_{\theta_{3}}(e_{{}_{G}},F_{e},z)=\pi_{\theta_{2}}(o(k)),$ (13)
which shows the left invariance of the G-CompACT policy on the end-effector frame. ∎
In what follows, we demonstrate that the pose signal produced by G-CompACT is equivariant when described in the spatial frame.
Corollary 1 ($SE(3)$ (Left) Equivariance of G-CompACT).
Let the output of G-CompACT in the spatial frame be $\bar{g}_{d}$ via (without considering the outputs related to gains)
$\bar{g}_{d}=g\!\cdot\!\pi_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w}).$ (14)
Then,
$g_{l}\bar{g}_{d}$ $=g_{l}g\cdot\pi_{\theta_{2}}(g_{l}\circ e_{{}_{G}}(g,g_{{}_{EDF}}),g_{l}\circ F_{e},g_{l}\circ I_{w})$
$=g_{l}g\cdot\pi_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w}).$ (15)
Since the temporary desired pose on the spatial frame signal, $\bar{g}_{d}$, is $SE(3)$ equivariant, the temporal aggregated signal $g_{d}$ is also equivariant. Therefore, by combining the G-CompACT with a spatial frame representation and temporal ensemble, we will now define the final form of G-CompACT as $\hat{\pi}_{\theta_{2}}$, which is given by:
$(g_{d},K_{p},K_{R})=\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w}).$ (16)
and satisfies the following equivariance property:
$\begin{split}&(g_{l}g_{d},K_{p},K_{R})(k)=\hat{\pi}_{\theta_{2}}(g_{l}\circ o(k)).\end{split}$ (17)
@@ IV-C EquiContact
The proposed EquiContact method comprises the high-level Diff-EDF, which serves as a high-level vision planner that provides the reference frame to be fed to GCEV, and the low-level G-CompACT, which handles fine-grained movement and contact interaction during insertion using real-time vision and force feedback. The overall pipeline of the EquiContact is presented in Fig. 1. Here, we demonstrate that the proposed EquiContact possesses the equivariance property. Let EquiContact be written as $h_{\theta}$ so that $h_{\theta}(g,g_{ref},F_{e})\mapsto f_{{}_{G}}$, i.e., $h_{\theta}:SE(3)\times SE(3)\times se^{*}(3)\to se^{*}(3)$.
@@
The EquiContact Policy $h_{\theta}$ is equivariant if it is described relative to the spatial frame.
Proof.
Suppose that the Assumption 1 and 2 hold. Let the object of interest, e.g., a peg for the picking task and a hole for the placing task, be observed by $\mathcal{O}^{ref}$, $I_{w}$ with its pose given by $g_{ref}$, so that the left-translated $g_{l}\cdot g_{ref}$ is observed by $g_{l}\circ\mathcal{O}^{ref}$ from the point cloud, and $g_{l}\circ I_{w}$ by the left-translated end-effector attached wrist camera as described in Fig. 3.
First, notice that $h_{\theta}$ can be fully written as
$h_{\theta}(g,g_{ref},F_{e})$ $=f_{{}_{G}}(g,\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g,f_{\theta_{1}}(\mathcal{O}^{ref})),F_{e},I_{w}))$ (18)
$=f_{{}_{G}}(g,\underset{=(g_{d},K_{p},K_{R})}{\underbrace{\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w})}}).$
Then, when both $g$ and $g_{ref}$ undergoes a left transformation $g_{l}$, from Assumption 1 and Corollary 1, the following holds:
$h_{\theta}(g_{l}g,g_{l}g_{ref},g_{l}\circ F_{e})$
$=f_{{}_{G}}(g_{l}g,\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g_{l}g,f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{ref})),g_{l}\circ F_{e},g_{l}\circ I_{w}))$
$=f_{{}_{G}}(g_{l}g,\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g_{l}g,g_{l}g_{{}_{EDF}}),g_{l}\circ F_{e},g_{l}\circ I_{w}))$
$=f_{{}_{G}}(g_{l}g,g_{l}g_{d},K_{p},K_{R})=f_{{}_{G}}(g,g_{d},K_{p},K_{R})$
$=h_{\theta}(g,g_{ref},F_{e}).$ (19)
We note that the second-last equation ($SE(3)$ left-invariance of the elastic wrench) comes from Lemma 1 of [11]. Finally, from the result of Proposition 2 of [11], it follows that
$h^{s}_{\theta}(g_{l}g,g_{l}g_{ref},g_{l}\circ F_{e})=\text{Ad}_{g_{l}^{\text{-}1}}^{T}h_{\theta}^{s}(g,g_{ref},F_{e}),$ (20)
where Ad is an Adjoint operator, $\text{Ad}_{g_{l}^{\text{-}1}}^{T}$ is a representation for $se^{*}(3)$ (wrench) domain for the group action $g_{l}$ [11], and superscript $s$ denotes a vector represented in the spatial frame. ∎
@@ V-A Benchmark Results
To validate the effectiveness of our proposed EquiContact framework and its underlying design principles, we compare EquiContact against three benchmark approaches: ACT with world-frame observation and action representation, executed with and without GAC, and CompACT [17]. We note that the GAC serves as a standard admittance control when the platform’s pose is fixed, i.e., untilted. The observation and action representation of the benchmarks and the proposed EquiContact are listed in Table I. We conducted a series of benchmark tests to demonstrate the effectiveness of each component in the proposed approach. The overall results are summarized in Table I.
@@ V-A1 Demonstration of Compliance
We begin by evaluating the role of compliance using the same ACT model architecture, executed with and without the Geometric Admittance Control (GAC). Results for this comparison are shown in the 1st and 2nd rows of Table I. Without GAC, the ACT model shows significantly lower success rates. The failure mode of the ACT w/o GAC involves collision: as the robot approaches the platform, excessive contact forces trigger safety shutdowns, preventing the task from completing. Due to the safety issue, we limited the number of trials without GAC to 10. This result demonstrates that the presence of compliant control is almost a decision factor between success and failure for contact-rich tasks.
@@ V-A2 Demonstration of Equivariance
Although the CompACT succeeds in insertion tasks in trained scenarios without excessive force exertion, it fails to generalize to spatially unseen configurations. This is expected, as its observation and action representations are defined in the global spatial frame, which neither guarantees nor encourages equivariance. The result of applying CompACT to the translationally unseen cases is shown in the Table I $3$rd row - Flat Platform (OOD). Note that the flat platform is only utilized, meaning it is only randomized translationally. We tested for $10$ cases and did not conduct more tests because it resulted in a $0\%$ success rate.
@@
From these experimental results, we draw the following key takeaway.
Remark 1 (Anchoring a Localized Policy on the Reference Frame for Equivariance).
The G-CompACT policy $\pi_{\theta}$ operates exclusively on localized inputs: GCEV, FT values in the end-effector frame, and images from wrist cameras. Its outputs, relative poses, and admittance gains are likewise defined in the local end-effector frame. The localized policy is anchored to a reference frame generated by the Diff-EDF planner. This architectural design induces spatial generalization and $SE(3)$-equivariance. Under this principle, one may hypothesize the reason for the spatial generalizability of recent vision-language-action models, which utilize both wrist and external cameras. The wrist camera provides a localized policy, anchored in the reference frame obtained from the external cameras and proprioceptive information. In fact, learning a reference frame from proprioceptive information (e.g., joint position) and the external camera is subtle because of its black-box nature. In contrast, our method, which utilizes a point cloud-based reference frame approach, is more interpretable. Thus, we propose “anchoring localized policies on globally estimated reference frames” as a general framework for the divide-and-conquer philosophy to enhance spatial generalization and interpretability. In our case, this structure also enables provable $SE(3)$ equivariance via differential geometric design.
V-B Failure Cases and Performance Recovery via Data Augmentation
@@
Although the EquiContact framework is vision-to-force equivariant in theory, this guarantee holds only under Assumptions 1 and 2. In particular, Assumption 2, which requires left-invariance of the visual features, is more challenging to enforce in practice, as our approach does not explicitly encode this property through loss functions or architectural inductive bias. As a result, the proposed G-CompACT algorithm shows degraded performance in scenarios with visual distractors or on a more severely tilted platform $(45^{\circ})$ – see Table III (“Base Dataset” column). Note that the larger tilting angle results in more out-of-distribution images, as the wrist cameras begin to capture more unseen background scenes, as illustrated in Fig. 3. To accommodate this, we augment the base dataset with $20$ demos collected using the visual distractor on a fixed, flat platform and $20$ demos collected from a $30^{\circ}$ angle of the tilted platform. After training with the augmented dataset, the success rates return to normal levels, as shown in the “Aug. Dataset” column of Table III, demonstrating the performance recovery achieved through data augmentation.
@@ VI Conclusion
In this work, we introduced EquiContact, a vision-to-force equivariant policy for spatially generalizable contact-rich tasks. By integrating a global reference frame estimator (Diff-EDF) with a fully localized visuomotor servoing policy module (G-CompACT), we demonstrate how compliance, localized policy, and induced equivariance can be unified to enable the peg-in-hole (PiH) task, a representative contact-rich precision task, under spatial perturbations. We proved the $SE(3)$ equivariance of the policy under assumptions on point cloud and image observations and validated its effectiveness through real-world experiments on PiH benchmarks. Compared to benchmark methods, our approach generalizes to unseen platform positions and orientations while maintaining low contact force and near-perfect success rates. Through extensive benchmark studies, we highlighted the effectiveness of the three principles – compliance, localized policy, and induced equivariance – for achieving spatial generalizability in contact-rich manipulation. We conclude that these principles offer a simple yet powerful design guideline for developing spatially generalizable and interpretable robotic policies complementing recent trends in end-to-end visuomotor learning and enabling a structured divide-and-conquer approach.
@@
SI Implementation Details
In this section, we provide the details of the hardware code implementation. Since every hardware system has a different implementation, our goal is to provide a comprehensive overview and a guide to our code distribution. We will use the exact names of the nodes to enable readers to easily understand our codebase.
SI-A Implementation Pipeline
@@
SI-B Diffusion-EDF
SI-B1 Implementation Heuristics
@@
SI-B2 Accuracy of the Diff-EDF
Here, we present detailed error values of the Diff-EDF evaluated on the training dataset. We note that we followed the implementation pipeline of the original Diff-EDF package, but just utilized the two pick models as described in Section SI-B1. The resulting root mean squared error (RMSE) values are shown in Table S1. The rotational errors are calculated via $R_{e}=R_{gt}R_{pick}^{T}$ and converted to $xyz$ euler angle representation, where $pick$ is replaced by $place$ for the placing task.
TABLE S1: RMSE error values of Diff-EDFs. The dimension of translational error is $\mathrm{mm}$ and the rotational error is $\mathrm{deg}$.
Pick Place
Translation $x$ ($\mathrm{mm}$) 16.76 4.981
Translation $y$ ($\mathrm{mm}$) 5.886 7.422
Translation $z$ ($\mathrm{mm}$) 10.79 5.236
Rotation $x$ ($\mathrm{deg}$) 19.88 13.74
Rotation $y$ ($\mathrm{deg}$) 40.55 18.99
Rotation $z$ ($\mathrm{deg}$) 106.7 91.41
As mentioned in Section I, the error in translation ranges from $5\mathrm{mm}$ to $15\mathrm{mm}$, and rotational error can range up to $40^{\circ}$. We note that the $z$-directional rotation error can be arbitrarily high in our case because our PiH task is symmetric.
SI-C Tips on applying Geometric Admittance Control
@@
SI-C2 Handling Force-Torque Sensor Data
@@
SI-C3 Implementation Tips
As mentioned in Section. IV, the detailed implementation method depends on the low-level servo controller type. Typical controllers provided by robot manufacturers are joint position control, joint velocity control, end-effector position control, and end-effector velocity control. If a torque controller is readily available, then directly implementing a geometric impedance controller (GIC) is recommended, but joint friction compensation needs to be handled appropriately. The end-effector velocity control can be straightforwardly applied from (IV-B1), by directly using $V_{d}^{b}(k)$ as a desired end-effector velocity command. One caveat here is the $V^{b}$, which is body-frame end-effector velocity. The common velocity provided by the manufacturer is defined on the Cartesian (Spatial) frame. If end-effector velocity is only offered in the Cartesian frame $V\in\mathbb{R}^{6}$, then
@@
SI-D Teleoperation Details
The architecture of our teleoperation system is explained to make it easier to reproduce and adapt our work. The system is constructed modularly, with the human-operator interface decoupled from the robot controller. This allows greater adaptation to different robotic platforms and control algorithms. Two ROS nodes make up the majority of the architecture: the $\mathtt{RobotNode}$ (named as $\mathtt{robot\_gac}$ in the code implementation), which processes proprioceptive information and executes a Geometric Admittance Controller to carry out operations, and the $\mathtt{Teleoperation}$ (named as $\mathtt{teleop\_ros}$ in the code implementation) node, which handles operator input.
@@
SI-E G-CompACT Training Details
@@
where $n_{p}\sim\mathcal{U}(-0.02,0.02)\in\mathbb{R}^{3}$ and $n_{r}\sim\mathcal{U}(-8^{\circ},8^{\circ})\in\mathbb{R}^{3}$, which denotes $2\mathrm{cm}$ of translational and $8^{\circ}$ (degree) of rotational error.

### v2

@@ IV-B1 G-CompACT
where $g=(p,R)\in SE(3)$ is a current end-effector pose, $g_{{}_{EDF}}=(p_{{}_{EDF}},R_{{}_{EDF}})$ is the reference frame predicted by Diff-EDF, and $(\cdot)^{\vee}$ denotes the vee-map, a mapping from $so(3)$ (Lie algebra of $SO(3)$) to $\mathbb{R}^{3}$. The physical meaning of GCEV is an error vector in the current end-effector frame between the current end-effector pose and the reference frame, considering the $SE(3)$ manifold structure, i.e., non-Euclidean. For the details of GCEV $e_{{}_{G}}$, we refer to [11, 12].
@@
Figure 3: Effects of the left group action $g_{l}$ to the end-effector pose $g$ and the reference frame $g_{ref}$, and to the wrists cameras $I_{w,1}$ and $I_{w,2}$. As the left group action is applied on the end-effector, the wrist cameras start to see backgrounds other than the optical table. $\{s\}$ denotes a spatial frame, i.e., robot base frame or world frame.
@@
The proof is presented in Appendix. -B. In what follows, we present an equivariant property of the pose signal produced by G-CompACT when described in the spatial frame.
Corollary 1 ($SE(3)$ Equivariance of G-CompACT).
The G-CompACT $\pi_{\theta_{2}}$ represented in the spatial frame satisfies the following equivariance property:
$\begin{split}&(g_{l}g_{d},K_{p},K_{R})(k)=\pi_{\theta_{2}}(g_{l}\circ o(k)).\end{split}$ (8)
@@
Notice that the desired pose signal in the spatial frame $g_{d}$ is obtained via $g_{d}=gg_{rel}=g\pi_{\theta_{2}}(o(k))$. Then, utilizing the left-invariance property, $g_{l}g\cdot\pi_{\theta_{2}}(g_{l}\circ o(k))=g_{l}g_{d}$. ∎
@@
IV-B2 Geometric Admittance Control (GAC)
We implement the geometric impedance control (GIC) proposed in [12, 13] in the geometric admittance control (GAC) setup [11]. Let the end-effector pose be denoted as $g\in SE(3)$ in a homogeneous matrix representation, or simply $g=(p,R)$, where $p\in\mathbb{R}^{3}$ is a position of the end-effector and $R\in SO(3)$ is a rotation matrix of the end-effector. The GAC operates with the $(g_{d},K_{p},K_{R})$ signal calculated from G-CompACT, where the desired end-effector pose is calculated via $g_{d}=gg_{rel}$. Given $g_{d}=(p_{d},R_{d})$, the desired end-effector dynamics for the GAC setup is written as follows:
$M\dot{V}^{b}+K_{d}V^{b}+f_{{}_{G}}=F_{e},$ (9)
where $M\in\mathbb{R}^{6\times 6}$ is symmetric positive definite desired inertia matrix, $K_{d}\in\mathbb{R}^{6\times 6}$ symmetric positive definite damping matrix, $F_{e}\in\mathbb{R}^{6}$ is external wrench applied to the end-effector in end-effector body frame and $V^{b}\in\mathbb{R}^{6}$ is a body-frame end-effector velocity. $K_{d}$ matrix is selected to ensure overdamped system, as $K_{d}=3\cdot\text{blkdiag}(\sqrt{K_{p}},\sqrt{K_{R}})$. Further, $f_{{}_{G}}=f_{{}_{G}}(g,g_{d},K_{p},K_{R})\in\mathbb{R}^{6}$ is a elastic wrench given by:
[…]
where $K_{p},K_{R}\in\mathbb{R}^{3\times 3}$ symmetric being positive stiffness matrices for the translational and rotational dynamics, respectively. Following the standard admittance control formulation, the desired end-effector pose command is calculated using (9), which is passed to the lowest level tracking controller. For details on GIC/GAC, we refer readers to [12, 11]. The admittance control loop is implemented at a $200\mathrm{Hz}$ rate using ROS2.
@@ IV-C EquiContact
The proposed EquiContact method comprises the high-level Diff-EDF, which serves as a high-level vision planner that provides the reference frame to be fed to GCEV, and the low-level G-CompACT, which handles fine-grained movement and contact interaction during insertion using real-time vision and force feedback. The overall pipeline of the EquiContact is presented in Fig. 1, and also summarized in Algorithm. 1. In Proposition 2, we demonstrate that the proposed EquiContact possesses the equivariance property. Let EquiContact be written as $h_{\theta}$ so that $h_{\theta}(g,g_{ref},F_{e})\mapsto f_{{}_{G}}$, i.e., $h_{\theta}:SE(3)\times SE(3)\times\mathbb{R}^{6}\to\mathbb{R}^{6}$.
@@
Suppose that the Assumption 1 and 2 hold. The EquiContact Policy $h_{\theta}$ is equivariant if it is described relative to the spatial frame.
The proof is shown in the Appendix. -B.
Algorithm 1 Inference Procedure of EquiContact
1:Diff-EDF $f_{\theta_{1}}$, G-CompACT $\pi_{\theta_{2}}$, Task $\in$ {pick, place}
2:Get scene and grasp point cloud $\mathcal{O}^{scene}$, $\mathcal{O}^{grasp}$
3:Run Diff-EDF for reference frame $g_{{}_{EDF}}\!=\!f_{\theta_{1}}(\mathcal{O}^{scene},\mathcal{O}^{grasp})$
4:Move the end-effector near the reference frame and
5:   initialize EquiContact $\pi_{\theta_{2}}$
6:for each inference timestep $k$ do
7:  Get current sensor values $g(k)$, $F_{e}(k)$, $I_{w}(k)$
8:  Calculate GCEV $e_{{}_{G}}(k)=e_{{}_{G}}(g(k),g_{{}_{EDF}})$ (4)
9:  Run G-CompACT:
10:      $(g_{rel},K_{p},K_{R})(k)=\pi_{\theta_{2}}(e_{{}_{G}},F_{e},I_{w})(k)$ (3)
11:  Calculate desired pose $g_{d}(k)=g(k)g_{rel}(k)$
12:  Update $(g_{d},K_{p},K_{R})(k)$ for GAC loop
13:  Run GAC realizing desired dynamics (9), (10)
14:end for
@@ V-A Benchmark Results
To evaluate the effectiveness of EquiContact and its core design principles, we compare it against three baselines: ACT with world-frame observations and actions, executed with and without GAC, and CompACT [17]. Note that GAC reduces to standard admittance control when the platform is fixed, i.e., not tilted. Table. I summarizes the observation/action representations used in each method and reports the benchmark results across all setups.
@@ V-A1 Demonstration of Compliance
As the importance of left-invariant compliant control action has already been verified in [11], we focus on verifying the necessity of compliant control action. We begin by evaluating the role of compliance using the same ACT model architecture, executed with and without the Geometric Admittance Control (GAC). Results for this comparison are shown in the 1st and 2nd rows of Table I. Without GAC, the ACT model shows significantly lower success rates. The failure mode of the ACT w/o GAC involves collision: as the robot approaches the platform, excessive contact forces trigger safety shutdowns, preventing the task from completing. Due to the safety issue, we limited the number of trials without GAC to 10. This result demonstrates that compliant control is nearly a deciding factor between success and failure in contact-rich tasks.
@@ V-A2 Demonstration of Equivariance
Although the CompACT succeeds in insertion tasks in trained scenarios without excessive force exertion, it fails to generalize to spatially unseen configurations. This is expected, as its observation and action representations are defined in the global spatial frame, which neither guarantees nor encourages equivariance. The result of applying CompACT to the translationally unseen cases is shown in the Table I 33rd row - Flat Platform (OOD). Note that the flat platform is only utilized, meaning it is only randomized translationally. We tested for $10$ cases and did not conduct more tests because it resulted in a $0\%$ success rate.
@@
V-B Limitations and Discussions
@@
V-B1 Limitation on Left-invariant Visual Features
Although EquiContact is theoretically vision-to-force equivariant, this guarantee holds only under Assumptions 1 and 2. In particular, Assumption 2, which requires left-invariant visual features, is challenging to ensure in practice, as our approach does not explicitly encode this property through loss functions or architectural inductive bias. Consequently, G-CompACT shows degraded performance in settings with visual distractors or severe tilting ($45^{\circ}$), as shown in Table III (“Base Dataset”). The degradation arises because larger tilt angles expose unseen background regions in the wrist camera view (see Fig. 3), resulting in out-of-distribution observations. This can be mitigated by augmenting the training data with visual distractors and demonstrations collected from a tilted angle. To demonstrate this, we added $20$ demos with distractors on a flat platform and 20 demos from a $30^{\circ}$ tilted setup. As shown in the “Aug. Dataset” column of Table III, the success rates have recovered to original levels.
Enforcing left-invariance on RGB images via architectural design remains an open challenge. Recent works have explored inducing 3D equivariance from 2D images [32, 33]. Another promising direction involves using pretrained foundation models, such as CLIP and DINO, which may exhibit implicit equivariance due to large-scale and diverse training. Language-guided features (e.g., using the prompt “peg”) can help extract relevant features across views. Alternatively, real-time segmentation models can be used to focus only on the object of interest. We will investigate enforcing left-invariance of images in future work.
V-B2 Hardware and Algorithmic Limitation
A key challenge in implementing equivariant policies for robotic manipulation is handling singularities [10]. When the robot is near singular, pose tracking accuracy degrades because controllers sacrifice tracking to avoid singularity, leading to poorly executed policy commands. This results in a distributional shift, which is a critical limitation for imitation learning. From our experimental demonstration perspective, a small workspace of the Diff-EDF (Fig. 2) was a drawback due to its limited receptive field. A combination of the singularity, joint limits, and small workspace resulted in limited demonstration of our approach, not covering a complete $SE(3)$ transformation.
V-C Validating Feasibility to Other Contact-rich Tasks
To further validate the EquiContact framework, we test it on two additional contact-rich tasks: screwing and surface wiping (see Fig. 5). In the screwing task, the robot aligns and screws a peg; in the wiping task, it erases a line from a whiteboard. For simplicity, we use a white background to eliminate visual distractions and assume that reference frames are known. The reference frame for screwing is the end-effector pose at full insertion; for wiping, it is the center of the board. As in the PiH setup, demonstrations are collected on a fixed platform and evaluated on out-of-distribution configurations, including tilted platforms. We collected 50 demonstrations per task. Results in Table IV show consistent success rates across all conditions, confirming EquiContact’s spatial generalizability.
Figure 5: (Left) Screwing task, (Right) Surface wiping (erasing) task. The same platform structure of the PiH is used.
TABLE IV: Success Rates of the G-CompACT for screwing and surface wiping tasks. The evaluation is conducted with the ground-truth reference frames. Both tasks are trained with $50$ demos.
Test Scenario Screwing Surface Wiping
Flat Platform (OOD) 10 / 10 10 / 10
Tilted Platform ($30^{\circ}$, OOD) 9 / 10 9 / 10
@@ VI Conclusion
In this work, we introduced EquiContact, a vision-to-force equivariant policy for spatially generalizable contact-rich tasks. By integrating a global reference frame estimator (Diff-EDF) with a fully localized visuomotor servoing policy module (G-CompACT), we demonstrate how compliance, localized policy, and induced equivariance can be unified to enable the peg-in-hole (PiH) task, a representative contact-rich precision task, under spatial perturbations. We proved the $SE(3)$ equivariance of the policy under assumptions on point cloud and image observations, validated its effectiveness through real-world experiments on PiH benchmarks and its feasibility towards screwing and surface wiping tasks. Compared to benchmark methods, our approach generalizes to unseen platform positions and orientations while maintaining low contact force and near-perfect success rates. Through extensive benchmark studies, we highlighted the effectiveness of the three principles – compliance, localized policy, and induced equivariance – for achieving spatial generalizability in contact-rich manipulation. We conclude that these principles offer a simple yet powerful design guideline for developing spatially generalizable and interpretable robotic policies complementing recent trends in end-to-end visuomotor learning and enabling a structured divide-and-conquer approach.
@@
Implementation Details
In this section, we provide the details of the hardware code implementation. Since every hardware system has a different implementation, our goal is to provide a comprehensive overview and a guide to our code distribution. We will use the exact names of the nodes to enable readers to understand our codebase easily. The hardware experimental implementation code is provided in https://github.com/Joohwan-Seo/EquiContact-Experiment, and the simulation implementation code in https://github.com/Joohwan-Seo/EquiContact-Simulation.
S-A Implementation Pipeline
@@
S-B Diffusion-EDF
S-B1 Implementation Heuristics
@@
S-C Tips on applying Geometric Admittance Control
@@
S-C2 Handling Force-Torque Sensor Data
@@
S-C3 Implementation Tips
As mentioned in Section. IV, the detailed implementation method depends on the low-level servo controller type. Typical controllers provided by robot manufacturers are joint position control, joint velocity control, end-effector position control, and end-effector velocity control. If a torque controller is readily available, then directly implementing a geometric impedance controller (GIC) is recommended, but joint friction compensation needs to be handled appropriately. Given the desired dynamics (9), the desired end-effector pose command $\tilde{g}_{d}(k)$ provided to the end-effector controller is calculated in discrete time as
$V_{d}^{b}(k)\!$ $=\!V^{b}(k)\!+\!T_{s}\cdot M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),$
$\tilde{g}_{d}(k)\!$ $=\!g(k)\cdot\exp{(\hat{V}_{d}^{b}(k)\cdot T_{s})},$ (16)
where $T_{s}$ is a sampling time ($5\mathrm{ms}$ in our case for GAC) and $\hat{(\cdot)}$ denotes a hat-map.
The end-effector velocity control can be straightforwardly applied from (S-C3), by directly using $V_{d}^{b}(k)$ as a desired end-effector velocity command. One caveat here is the $V^{b}$, which is body-frame end-effector velocity. The common velocity provided by the manufacturer is defined on the Cartesian (Spatial) frame. If end-effector velocity is only offered in the Cartesian frame $V\in\mathbb{R}^{6}$, then
@@
S-D Teleoperation Details
The architecture of our teleoperation system is explained to make it easier to reproduce and adapt our work. The system is constructed modularly, with the human-operator interface decoupled from the robot controller. This allows greater adaptation to different robotic platforms and control algorithms. The teleoperation function is implemented by two ROS nodes: the $\mathtt{RobotNode}$ (named as $\mathtt{robot\_gac}$ in the code implementation), which processes proprioceptive information and executes a Geometric Admittance Controller to carry out operations, and the $\mathtt{Teleoperation}$ (named as $\mathtt{teleop\_ros}$ in the code implementation) node, which handles operator input.
@@
S-E G-CompACT Training Details
@@
where $n_{p}\sim\mathcal{U}(-0.02,0.02)\in\mathbb{R}^{3}$ and $n_{r}\sim\mathcal{U}(-8^{\circ},8^{\circ})\in\mathbb{R}^{3}$, which denotes $2\mathrm{cm}$ of translational and $8^{\circ}$ (degree) of rotational error. The impact of adding noise to the reference frame is further shown in Appendix. -C.